OpenStudy (anonymous):

i never understood this question, If the slope of a straight line is 0 and the y-intercept is –2, what is the equation of the line?

OpenStudy (ivettef365):

the formula for the equation is y = mx + b, so just plug in numbers

OpenStudy (ivettef365):

m = slope = 0 b = y-int = -2 so you have y = (0)x -2 so it is y = -2

OpenStudy (anonymous):

another question

OpenStudy (anonymous):

What is the slope of the line 8x + 6y = 12?

OpenStudy (ivettef365):

again, if you convert to the formula y = mx + b, by isolating the y

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

or 3/4?

OpenStudy (ivettef365):

so you have 8x + 6y = 12 6y = -8x + 12 y = -8/6x +12/6 so y = -4/3x +2

OpenStudy (ivettef365):

so what number is in the position of m

OpenStudy (anonymous):

-4/3 got cha, thanks a lot

OpenStudy (ivettef365):

correct

OpenStudy (ivettef365):

yw :)